# Geometric calculations on probability manifolds from reciprocal relations in Master equations

### v4

@@ 1. Introductions
Non-equilibrium thermodynamics [5, 8] studies dynamical systems that interact with reservoirs for thermal energy or volume. Examples are coupled physical irreversible processes in complex systems, which are modeled by probability functions over physical states. Typical dynamics include thermoelectric phenomena and heat conduction in an anisotropic medium. In the modeling of irreversible processes, Onsager introduces the reciprocal relationship [22], which describes the symmetric dissipation of the probability transition equations arising from physical processes, known as free-energy dissipation or entropy production. Here, the free energy dissipation follows the second law of thermodynamics. Rather than addressing arbitrary physical mechanisms, one can focus on two canonical classes of stochastic models: overdamped Langevin dynamics in continuous state spaces, and detailed-balance Markov processes with transition rates on discrete domains. In both cases, entropy production is expressed as the time derivative of an entropy (or free-energy) functional along the governing probability evolution equations, such as the Fokker–Planck equation in the continuous setting or the master equation in the discrete setting.
@@
In this paper, we study Riemannian calculations in a class of thermodynamical probability manifolds. We first compute the Levi-Civita connection, gradient, and Hessian operators on probability manifolds. We further derive Riemannian and sectional curvatures in the probability simplex set. Two examples of these computations are performed. One example is the Levi-Civita connection on a three-state probability simplex, where the discrete probability manifold is constructed from the master equation for a chemical monomolecular triangle reaction. The other example is the Ricci curvature on a probability manifold with a three-point lattice graph.
@@ 2. Onsager reciprocal relations, network notations, and Thermodynamical probability manifolds
This section reviews finite-state master equations with detailed balance relations, Onsager reciprocal relations, network notation, and a class of thermodynamical probability manifolds. These studies review results in [20, 21, 25].
@@ 2.2. Onsager reciprocal relations
In fact, the R.H.S. of symmetric formulation (4) can be written as a matrix function times a vector function. This is known as the strong Onsager gradient flow (physical terminology), which contains a symmetric matrix function (Onsager’s response matrix) and the gradient of a free-energy function (the force). To illustrate this point, we first define a force vector function.
@@ 3.1. Levi-Civita connection, parallel transport and geodesics
    In detail,
    $\left\{\begin{split}&\frac{d\gamma_{i}}{dt}+\mathrm{div}_{\omega}(\theta(\gamma)\nabla_{\omega}\Phi)_{i}=0,\\ &\frac{d\Phi_{i}}{dt}+\frac{1}{2}\sum_{j\in N(i)}(\nabla_{\omega}\Phi)^{2}_{ij}\partial_{i}\theta_{ij}(\gamma)=c(t),\end{split}\right$.
    where $c\colon[0,T]\rightarrow\mathbb{R}$ is a scalar function of the time variable $t$.
@@
We note that $L(\theta)u_{0}=0$, where $u_{0}=\frac{1}{\sqrt{n}}(1,\cdots,1)^{\mathsf{T}}\in\mathbb{R}^{n}$ is a constant vector. When $\gamma\in\mathcal{P}_{+}$, $u_{0}$ is the only basis for the kernel space of $L(\theta)$. Thus, there exists a constant vector $c(t)(1,\cdots,1)^{\mathsf{T}}$, such that
@@
This finishes the proof. ∎
@@ 3.2. Riemannian curvature tensor
$\begin{split}\langle\bar{\mathrm{R}}(\mathsf{V}_{\Phi_{1}},\mathsf{V}_{\Phi_{2}})\mathsf{V}_{\Phi_{3}},\mathsf{V}_{\Phi_{4}}\rangle=&\langle\bar{\nabla}_{\mathsf{V}_{\Phi_{1}}}\bar{\nabla}_{\mathsf{V}_{\Phi_{2}}}\mathsf{V}_{\Phi_{3}}-\bar{\nabla}_{\mathsf{V}_{\Phi_{2}}}\bar{\nabla}_{\mathsf{V}_{\Phi_{1}}}\bar{\mathsf{V}}_{\Phi_{3}}-\bar{\nabla}_{[\mathsf{V}_{\Phi_{1}},\mathsf{V}_{\Phi_{2}}]}\mathsf{V}_{\Phi_{3}},\mathsf{V}_{\Phi_{4}}\rangle\\ =&\quad\mathsf{V}_{\Phi_{1}}\langle\bar{\nabla}_{\mathsf{V}_{\Phi_{2}}}\mathsf{V}_{\Phi_{3}},\mathsf{V}_{\Phi_{4}}\rangle-\langle\bar{\nabla}_{\mathsf{V}_{\Phi_{2}}}\mathsf{V}_{\Phi_{3}},\bar{\nabla}_{\mathsf{V}_{\Phi_{1}}}\mathsf{V}_{\Phi_{4}}\rangle\\ &-\mathsf{V}_{\Phi_{2}}\langle\nabla_{\mathsf{V}_{\Phi_{1}}}\mathsf{V}_{\Phi_{3}},\mathsf{V}_{\Phi_{4}}\rangle+\langle\nabla_{\mathsf{V}_{\Phi_{1}}}\mathsf{V}_{\Phi_{3}},\bar{\nabla}_{\mathsf{V}_{\Phi_{2}}}\mathsf{V}_{\Phi_{4}}\rangle\\ &-\langle\nabla_{[\mathsf{V}_{\Phi_{1}},\mathsf{V}_{\Phi_{2}}]}\mathsf{V}_{\Phi_{3}},\mathsf{V}_{\Phi_{4}}\rangle.\end{split}$ (28)
@@
$\begin{split}&\langle\nabla_{[\mathsf{V}_{\Phi_{1}},\mathsf{V}_{\Phi_{2}}]}\mathsf{V}_{\Phi_{3}},\mathsf{V}_{\Phi_{4}}\rangle\\ =&\frac{1}{2}\Phi_{4}^{\mathsf{T}}\Big\{L(\mathsf{V}_{\Phi}\theta)\Phi_{3}-L(\mathsf{V}_{\Phi_{3}}\theta)\Phi+L(\theta)\Gamma(\Phi,\Phi_{3},p)\Big\}\\ =&\frac{1}{2}\Phi_{4}^{\mathsf{T}}L(\mathsf{V}_{\Phi}\theta)\Phi_{3}-\frac{1}{2}\Phi_{4}^{\mathsf{T}}L(\mathsf{V}_{\Phi_{3}}\theta)\Phi+\frac{1}{2}\Phi_{3}^{\mathsf{T}}L(\mathsf{V}_{\Phi_{4}}\theta)\Phi\\ =&\frac{1}{2}\Phi_{4}^{\mathsf{T}}L([\mathsf{V}_{\Phi_{1}},\mathsf{V}_{\Phi_{2}}]\theta)\Phi_{3}+\frac{1}{2}[\mathsf{V}_{\Phi_{4}},\mathsf{V}_{\Phi_{3}}]^{\mathsf{T}}R(\theta)[\mathsf{V}_{\Phi_{1}},\mathsf{V}_{\Phi_{2}}]\\ =&\frac{1}{2}\Phi_{4}^{\mathsf{T}}L(\mathsf{V}_{\Phi_{1}}\mathsf{V}_{\Phi_{2}}\theta)\Phi_{3}-\frac{1}{2}\Phi_{4}^{\mathsf{T}}L(\mathsf{V}_{\Phi_{2}}\mathsf{V}_{\Phi_{1}}\theta)\Phi_{3}+\frac{1}{2}[\mathsf{V}_{\Phi_{4}},\mathsf{V}_{\Phi_{3}}]^{\mathsf{T}}R(\theta)[\mathsf{V}_{\Phi_{1}},\mathsf{V}_{\Phi_{2}}].\end{split}$ (33)
@@
We finish the proof of equation (27). ∎
We present an explicit formulation of the Riemannian curvature $\bar{\mathrm{R}}$. Denote a third order iterative Gamma tensor $\Gamma^{3}\colon\mathbb{R}^{n}/\mathbb{R}\times\mathbb{R}^{n}/\mathbb{R}\times\mathbb{R}^{n}/\mathbb{R}\times\mathbb{R}^{n}/\mathbb{R}\times\mathcal{P}_{+}\rightarrow\mathbb{R}^{n\times n}$. Given vectors $\Phi_{1}$, $\Phi_{2}$, $\Phi_{3}$, $\Phi_{4}\in\mathbb{R}^{n}$, write $\Gamma^{3}(\Phi_{1},\Phi_{2},\Phi_{3},\Phi_{4},p)=(\Gamma^{3}(\Phi_{1},\Phi_{2},\Phi_{3},\Phi_{4},p)_{ij})_{1\leq i,j\leq n}$, such that
$(\Gamma^{3}(\Phi_{1},\Phi_{2},\Phi_{3},\Phi_{4},p))_{ij}:=\frac{1}{2}\sum_{k=1}^{n}(\nabla_{\omega})_{ik}\Big((\nabla_{\omega}\Gamma(\Phi_{1},\Phi_{2},p))_{ij}(\nabla_{\omega}\Phi_{4})_{ij}\frac{\partial\theta_{ij}}{\partial p_{i}}\Big)(\nabla_{\omega}\Phi_{3})_{ik}\theta_{ik}.$
We also denote $C=(C_{ij})_{1\leq i,j\leq n}\in\mathbb{R}^{n\times n}$, such that
@@
$\begin{split}&\langle\bar{\mathrm{R}}(\mathsf{V}_{\Phi_{1}},\mathsf{V}_{\Phi_{2}})\mathsf{V}_{\Phi_{3}},\mathsf{V}_{\Phi_{4}}\rangle\\ =&\quad\frac{1}{2}\sum_{i,j,k,l=1}^{n}\frac{\partial^{2}\theta_{ij}}{\partial p_{i}\partial p_{i}}\theta_{ik}\theta_{il}\sqrt{\omega_{ik}}\sqrt{\omega_{il}}\Big\{-(\nabla_{\omega}\Phi_{2})_{ij}(\nabla_{\omega}\Phi_{4})_{ij}(\nabla_{\omega}\Phi_{1})_{ik}(\nabla_{\omega}\Phi_{3})_{il}\\ &\hskip 165.02606pt-(\nabla_{\omega}\Phi_{1})_{ij}(\nabla_{\omega}\Phi_{3})_{ij}(\nabla_{\omega}\Phi_{2})_{ik}(\nabla_{\omega}\Phi_{4})_{il}\\ &\hskip 165.02606pt+(\nabla_{\omega}\Phi_{2})_{ij}(\nabla_{\omega}\Phi_{3})_{ij}(\nabla_{\omega}\Phi_{1})_{ik}(\nabla_{\omega}\Phi_{4})_{il}\\ &\hskip 165.02606pt+(\nabla_{\omega}\Phi_{1})_{ij}(\nabla_{\omega}\Phi_{4})_{ij}(\nabla_{\omega}\Phi_{2})_{ik}(\nabla_{\omega}\Phi_{3})_{il}\Big\}\\ &+\frac{1}{8}\sum_{i,j,k,l=1}^{n}\theta_{ij}\theta_{kl}\Big\{-(\nabla_{\omega})_{ij}(\nabla_{\omega})_{kl}(\nabla_{\omega}\Phi_{2}\nabla_{\omega}\Phi_{4}\nabla_{pp}^{2}\theta)(\nabla_{\omega}\Phi_{1})_{ij}(\nabla_{\omega}\Phi_{3})_{kl}\\ &\hskip 88.2037pt-(\nabla_{\omega})_{ij}(\nabla_{\omega})_{kl}(\nabla_{\omega}\Phi_{1}\nabla_{\omega}\Phi_{3}\nabla_{pp}^{2}\theta)(\nabla_{\omega}\Phi_{2})_{ij}(\nabla_{\omega}\Phi_{4})_{kl}\\ &\hskip 88.2037pt+(\nabla_{\omega})_{ij}(\nabla_{\omega})_{kl}(\nabla_{\omega}\Phi_{2}\nabla_{\omega}\Phi_{3}\nabla_{pp}^{2}\theta)(\nabla_{\omega}\Phi_{1})_{ij}(\nabla_{\omega}\Phi_{4})_{kl}\\ &\hskip 88.2037pt+(\nabla_{\omega})_{ij}(\nabla_{\omega})_{kl}(\nabla_{\omega}\Phi_{1}\nabla_{\omega}\Phi_{4}\nabla_{pp}^{2}\theta)(\nabla_{\omega}\Phi_{2})_{ij}(\nabla_{\omega}\Phi_{3})_{kl}\Big\}\\ &+\frac{1}{8}\sum_{(i,j)\in E}\theta_{ij}\Big\{-\Gamma^{3}(\Phi_{2},\Phi_{4},\Phi_{1},\Phi_{3},p)_{ij}-\Gamma^{3}(\Phi_{2},\Phi_{4},\Phi_{3},\Phi_{1},p)_{ij}\\ &\hskip 71.13188pt-\Gamma^{3}(\Phi_{1},\Phi_{3},\Phi_{2},\Phi_{4},p)_{ij}-\Gamma^{3}(\Phi_{1},\Phi_{3},\Phi_{4},\Phi_{2},p)_{ij}\\ &\hskip 71.13188pt+\Gamma^{3}(\Phi_{2},\Phi_{3},\Phi_{1},\Phi_{4},p)_{ij}+\Gamma^{3}(\Phi_{2},\Phi_{3},\Phi_{4},\Phi_{1},p)_{ij}\\ &\hskip 71.13188pt+\Gamma^{3}(\Phi_{1},\Phi_{4},\Phi_{2},\Phi_{3},p)_{ij}+\Gamma^{3}(\Phi_{1},\Phi_{4},\Phi_{3},\Phi_{2},p)_{ij}\\ &\hskip 56.9055pt+(\nabla_{\omega}\Gamma(\Phi_{1},\Phi_{3},p))_{ij}(\nabla_{\omega}\Gamma(\Phi_{2},\Phi_{4},p))_{ij}-(\nabla_{\omega}\Gamma(\Phi_{2},\Phi_{3},p))_{ij}(\nabla_{\omega}\Gamma(\Phi_{1},\Phi_{4},p))_{ij}\Big\}\\ &+\frac{1}{4}\Big\{[\mathsf{V}_{\Phi_{1}},\mathsf{V}_{\Phi_{3}}]^{\mathsf{T}}R(\theta)[\mathsf{V}_{\Phi_{2}},\mathsf{V}_{\Phi_{4}}]-[\mathsf{V}_{\Phi_{2}},\mathsf{V}_{\Phi_{3}}]^{\mathsf{T}}R(\theta)[\mathsf{V}_{\Phi_{1}},\mathsf{V}_{\Phi_{4}}]+2[\mathsf{V}_{\Phi_{3}},\mathsf{V}_{\Phi_{4}}]^{\mathsf{T}}R(\theta)[\mathsf{V}_{\Phi_{1}},\mathsf{V}_{\Phi_{2}}]\Big\}.\end{split}$ (36)

### v5

# arXiv:2504.19368v5[math-ph] 27 Feb 2026

@@ 1. Introductions
Non-equilibrium thermodynamics [5, 8] studies dynamical systems that interact with reservoirs for thermal energy or volume. Examples are coupled physical irreversible processes in complex systems, which are modeled by probability functions over physical states. Typical dynamics include thermoelectric phenomena and heat conduction in an anisotropic medium. In the modeling of irreversible processes, Onsager introduces the reciprocal relationship [22], which describes the symmetric dissipation of the probability transition equations arising from physical processes, known as free energy dissipation or entropy production. Here, the free energy dissipation follows the second law of thermodynamics. Rather than addressing arbitrary physical mechanisms, one can focus on two canonical classes of stochastic models: overdamped Langevin dynamics in continuous state spaces, and detailed-balance Markov processes with transition rates on discrete domains. In both cases, entropy production is expressed as the time derivative of an entropy (or free energy) functional along the governing probability evolution equations, such as the Fokker–Planck equation in the continuous setting or the master equation in the discrete setting.
@@
In this paper, we study Riemannian calculations in a class of thermodynamical probability manifolds. This study is a generalization of Otto calculus on discrete states with general mobility functions. We first compute the Levi-Civita connection, gradient, and Hessian operators on probability manifolds. We further derive Riemannian and sectional curvatures in the probability simplex set. Two examples of these computations are performed. One example is the Levi-Civita connection on a three-state probability simplex, where the discrete probability manifold is constructed from the master equation for a chemical monomolecular triangle reaction. The other example is the Ricci curvature on a probability manifold with a three-point lattice graph.
@@ 2. Onsager reciprocal relations, network notations, and Thermodynamical probability manifolds
This section reviews finite-state master equations with detailed balance relations, Onsager reciprocal relations, network notation, and a class of thermodynamical probability manifolds. See more details in [20, 21, 25].
@@ 2.2. Onsager reciprocal relations
In fact, the R.H.S. of symmetric formulation (4) can be written as a matrix function times a vector function. This is known as the strong Onsager gradient flow (physical terminology), which contains a symmetric matrix function (Onsager’s response matrix) and the gradient of a free energy function (the force). To illustrate this point, we first define a force vector function.
@@ 3.1. Levi-Civita connection, parallel transport and geodesics
    In detail, a solution of geodesic equation satisfies a system of continuity equation and Hamilton-Jacobi equation on graphs:
    $\left\{\begin{split}&\frac{d\gamma_{i}}{dt}+\mathrm{div}_{\omega}(\theta(\gamma)\nabla_{\omega}\Phi)_{i}=0,\\ &\frac{d\Phi_{i}}{dt}+\frac{1}{2}\sum_{j\in N(i)}(\nabla_{\omega}\Phi)^{2}_{ij}\partial_{i}\theta_{ij}(\gamma)=0.\end{split}\right$. (24)
@@
We note that $L(\theta)u_{0}=0$, where $u_{0}=\frac{1}{\sqrt{n}}(1,\cdots,1)^{\mathsf{T}}\in\mathbb{R}^{n}$ is a constant vector. When $\gamma\in\mathcal{P}_{+}$, $u_{0}$ is the only basis for the kernel space of $L(\theta)$. Thus, there exists a constant vector $c(t)(1,\cdots,1)^{\mathsf{T}}$, where $c\colon[0,T]\rightarrow\mathbb{R}$ is a scalar function of the time variable $t$, such that
@@
Assume $c(t)$ is integrable in the time interval $[0,T]$, and we let $\hat{\Phi}_{i}(t)=\Phi_{i}(t)+\int_{0}^{t}c(s)ds$, for any $i\in\{1,\cdots,n\}$, then $(\gamma(t),\hat{\Phi}(t))$ satisfies the geodesic equation (24). This finishes the proof. ∎
@@ 3.2. Riemannian curvature tensor
$\begin{split}\langle\bar{\mathrm{R}}(\mathsf{V}_{\Phi_{1}},\mathsf{V}_{\Phi_{2}})\mathsf{V}_{\Phi_{3}},\mathsf{V}_{\Phi_{4}}\rangle=&\langle\bar{\nabla}_{\mathsf{V}_{\Phi_{1}}}\bar{\nabla}_{\mathsf{V}_{\Phi_{2}}}\mathsf{V}_{\Phi_{3}}-\bar{\nabla}_{\mathsf{V}_{\Phi_{2}}}\bar{\nabla}_{\mathsf{V}_{\Phi_{1}}}\bar{\mathsf{V}}_{\Phi_{3}}-\bar{\nabla}_{[\mathsf{V}_{\Phi_{1}},\mathsf{V}_{\Phi_{2}}]}\mathsf{V}_{\Phi_{3}},\mathsf{V}_{\Phi_{4}}\rangle\\ =&\quad\mathsf{V}_{\Phi_{1}}\langle\bar{\nabla}_{\mathsf{V}_{\Phi_{2}}}\mathsf{V}_{\Phi_{3}},\mathsf{V}_{\Phi_{4}}\rangle-\langle\bar{\nabla}_{\mathsf{V}_{\Phi_{2}}}\mathsf{V}_{\Phi_{3}},\bar{\nabla}_{\mathsf{V}_{\Phi_{1}}}\mathsf{V}_{\Phi_{4}}\rangle\\ &-\mathsf{V}_{\Phi_{2}}\langle\bar{\nabla}_{\mathsf{V}_{\Phi_{1}}}\mathsf{V}_{\Phi_{3}},\mathsf{V}_{\Phi_{4}}\rangle+\langle\bar{\nabla}_{\mathsf{V}_{\Phi_{1}}}\mathsf{V}_{\Phi_{3}},\bar{\nabla}_{\mathsf{V}_{\Phi_{2}}}\mathsf{V}_{\Phi_{4}}\rangle\\ &-\langle\bar{\nabla}_{[\mathsf{V}_{\Phi_{1}},\mathsf{V}_{\Phi_{2}}]}\mathsf{V}_{\Phi_{3}},\mathsf{V}_{\Phi_{4}}\rangle.\end{split}$ (29)
@@
$\begin{split}&\langle\bar{\nabla}_{[\mathsf{V}_{\Phi_{1}},\mathsf{V}_{\Phi_{2}}]}\mathsf{V}_{\Phi_{3}},\mathsf{V}_{\Phi_{4}}\rangle\\ =&\frac{1}{2}\Phi_{4}^{\mathsf{T}}\Big\{L(\mathsf{V}_{\Phi}\theta)\Phi_{3}-L(\mathsf{V}_{\Phi_{3}}\theta)\Phi+L(\theta)\Gamma(\Phi,\Phi_{3},p)\Big\}\\ =&\frac{1}{2}\Phi_{4}^{\mathsf{T}}L(\mathsf{V}_{\Phi}\theta)\Phi_{3}-\frac{1}{2}\Phi_{4}^{\mathsf{T}}L(\mathsf{V}_{\Phi_{3}}\theta)\Phi+\frac{1}{2}\Phi_{3}^{\mathsf{T}}L(\mathsf{V}_{\Phi_{4}}\theta)\Phi\\ =&\frac{1}{2}\Phi_{4}^{\mathsf{T}}L([\mathsf{V}_{\Phi_{1}},\mathsf{V}_{\Phi_{2}}]\theta)\Phi_{3}+\frac{1}{2}[\mathsf{V}_{\Phi_{4}},\mathsf{V}_{\Phi_{3}}]^{\mathsf{T}}R(\theta)[\mathsf{V}_{\Phi_{1}},\mathsf{V}_{\Phi_{2}}]\\ =&\frac{1}{2}\Phi_{4}^{\mathsf{T}}L(\mathsf{V}_{\Phi_{1}}\mathsf{V}_{\Phi_{2}}\theta)\Phi_{3}-\frac{1}{2}\Phi_{4}^{\mathsf{T}}L(\mathsf{V}_{\Phi_{2}}\mathsf{V}_{\Phi_{1}}\theta)\Phi_{3}+\frac{1}{2}[\mathsf{V}_{\Phi_{4}},\mathsf{V}_{\Phi_{3}}]^{\mathsf{T}}R(\theta)[\mathsf{V}_{\Phi_{1}},\mathsf{V}_{\Phi_{2}}].\end{split}$ (34)
@@
We finish the derivation of equation (28). ∎
We present an explicit formulation of the Riemannian curvature $\bar{\mathrm{R}}$. Denote a third order iterative Gamma operator $\Gamma^{3}\colon\mathbb{R}^{n}/\mathbb{R}\times\mathbb{R}^{n}/\mathbb{R}\times\mathbb{R}^{n}/\mathbb{R}\times\mathbb{R}^{n}/\mathbb{R}\times\mathcal{P}_{+}\rightarrow\mathbb{R}^{n\times n}$. Given vectors $\Phi_{1}$, $\Phi_{2}$, $\Phi_{3}$, $\Phi_{4}\in\mathbb{R}^{n}$, write $\Gamma^{3}(\Phi_{1},\Phi_{2},\Phi_{3},\Phi_{4},p)=(\Gamma^{3}(\Phi_{1},\Phi_{2},\Phi_{3},\Phi_{4},p)_{ij})_{1\leq i,j\leq n}$, such that
$\begin{split}\Gamma^{3}(\Phi_{1},\Phi_{2},\Phi_{3},\Phi_{4},p)_{ij}:=&\frac{1}{2}\sum_{k=1}^{n}(\nabla_{\omega})_{ik}\Big((\nabla_{\omega}\Gamma(\Phi_{1},\Phi_{2},p))_{ij}(\nabla_{\omega}\Phi_{4})_{ij}\frac{\partial\theta_{ij}}{\partial p_{i}}\Big)(\nabla_{\omega}\Phi_{3})_{ik}\theta_{ik}.\end{split}$
Here, we denote a matrix $A=(A_{ij})_{i\leq i,j\leq n}\in\mathbb{R}^{n\times n}$, with $A_{ij}:=(\nabla_{\omega}\Gamma(\Phi_{1},\Phi_{2},p))_{ij}(\nabla_{\omega}\Phi_{4})_{ij}\frac{\partial\theta_{ij}}{\partial p_{i}}$, and
$(\nabla_{\omega})_{ik}A_{ij}:=\sqrt{\omega_{ik}}\Big(A_{kj}-A_{ij}\Big).$
We also denote a matrix $C=(C_{ij})_{1\leq i,j\leq n}\in\mathbb{R}^{n\times n}$, such that
@@
$\begin{split}&\langle\bar{\mathrm{R}}(\mathsf{V}_{\Phi_{1}},\mathsf{V}_{\Phi_{2}})\mathsf{V}_{\Phi_{3}},\mathsf{V}_{\Phi_{4}}\rangle\\ =&\quad\frac{1}{2}\sum_{i,j,k,l=1}^{n}\frac{\partial^{2}\theta_{ij}}{\partial p_{i}\partial p_{i}}\theta_{ik}\theta_{il}\sqrt{\omega_{ik}}\sqrt{\omega_{il}}\Big\{-(\nabla_{\omega}\Phi_{2})_{ij}(\nabla_{\omega}\Phi_{4})_{ij}(\nabla_{\omega}\Phi_{1})_{ik}(\nabla_{\omega}\Phi_{3})_{il}\\ &\hskip 165.02606pt-(\nabla_{\omega}\Phi_{1})_{ij}(\nabla_{\omega}\Phi_{3})_{ij}(\nabla_{\omega}\Phi_{2})_{ik}(\nabla_{\omega}\Phi_{4})_{il}\\ &\hskip 165.02606pt+(\nabla_{\omega}\Phi_{2})_{ij}(\nabla_{\omega}\Phi_{3})_{ij}(\nabla_{\omega}\Phi_{1})_{ik}(\nabla_{\omega}\Phi_{4})_{il}\\ &\hskip 165.02606pt+(\nabla_{\omega}\Phi_{1})_{ij}(\nabla_{\omega}\Phi_{4})_{ij}(\nabla_{\omega}\Phi_{2})_{ik}(\nabla_{\omega}\Phi_{3})_{il}\Big\}\\ &+\frac{1}{8}\sum_{i,j,k,l=1}^{n}\theta_{ij}\theta_{kl}\Big\{-(\nabla_{\omega})_{ij}(\nabla_{\omega})_{kl}(\nabla_{\omega}\Phi_{2}\nabla_{\omega}\Phi_{4}\nabla_{pp}^{2}\theta)(\nabla_{\omega}\Phi_{1})_{ij}(\nabla_{\omega}\Phi_{3})_{kl}\\ &\hskip 88.2037pt-(\nabla_{\omega})_{ij}(\nabla_{\omega})_{kl}(\nabla_{\omega}\Phi_{1}\nabla_{\omega}\Phi_{3}\nabla_{pp}^{2}\theta)(\nabla_{\omega}\Phi_{2})_{ij}(\nabla_{\omega}\Phi_{4})_{kl}\\ &\hskip 88.2037pt+(\nabla_{\omega})_{ij}(\nabla_{\omega})_{kl}(\nabla_{\omega}\Phi_{2}\nabla_{\omega}\Phi_{3}\nabla_{pp}^{2}\theta)(\nabla_{\omega}\Phi_{1})_{ij}(\nabla_{\omega}\Phi_{4})_{kl}\\ &\hskip 88.2037pt+(\nabla_{\omega})_{ij}(\nabla_{\omega})_{kl}(\nabla_{\omega}\Phi_{1}\nabla_{\omega}\Phi_{4}\nabla_{pp}^{2}\theta)(\nabla_{\omega}\Phi_{2})_{ij}(\nabla_{\omega}\Phi_{3})_{kl}\Big\}\\ &+\frac{1}{8}\sum_{(i,j)\in E}\Big\{\quad-\Gamma^{3}(\Phi_{2},\Phi_{4},\Phi_{1},\Phi_{3},p)_{ij}-\Gamma^{3}(\Phi_{2},\Phi_{4},\Phi_{3},\Phi_{1},p)_{ij}\\ &\hskip 71.13188pt-\Gamma^{3}(\Phi_{1},\Phi_{3},\Phi_{2},\Phi_{4},p)_{ij}-\Gamma^{3}(\Phi_{1},\Phi_{3},\Phi_{4},\Phi_{2},p)_{ij}\\ &\hskip 71.13188pt+\Gamma^{3}(\Phi_{2},\Phi_{3},\Phi_{1},\Phi_{4},p)_{ij}+\Gamma^{3}(\Phi_{2},\Phi_{3},\Phi_{4},\Phi_{1},p)_{ij}\\ &\hskip 71.13188pt+\Gamma^{3}(\Phi_{1},\Phi_{4},\Phi_{2},\Phi_{3},p)_{ij}+\Gamma^{3}(\Phi_{1},\Phi_{4},\Phi_{3},\Phi_{2},p)_{ij}\\ &\hskip 28.45274pt+\theta_{ij}\Big[(\nabla_{\omega}\Gamma(\Phi_{1},\Phi_{3},p))_{ij}(\nabla_{\omega}\Gamma(\Phi_{2},\Phi_{4},p))_{ij}-(\nabla_{\omega}\Gamma(\Phi_{2},\Phi_{3},p))_{ij}(\nabla_{\omega}\Gamma(\Phi_{1},\Phi_{4},p))_{ij}\Big]\Big\}\\ &+\frac{1}{4}\Big\{[\mathsf{V}_{\Phi_{1}},\mathsf{V}_{\Phi_{3}}]^{\mathsf{T}}R(\theta)[\mathsf{V}_{\Phi_{2}},\mathsf{V}_{\Phi_{4}}]-[\mathsf{V}_{\Phi_{2}},\mathsf{V}_{\Phi_{3}}]^{\mathsf{T}}R(\theta)[\mathsf{V}_{\Phi_{1}},\mathsf{V}_{\Phi_{4}}]+2[\mathsf{V}_{\Phi_{3}},\mathsf{V}_{\Phi_{4}}]^{\mathsf{T}}R(\theta)[\mathsf{V}_{\Phi_{1}},\mathsf{V}_{\Phi_{2}}]\Big\}.\end{split}$ (37)